Fermi National Laboratory


Accelerator Update for September 21, 2005 - September 23, 2005

Wednesday September 21
The day shift began with the Tevatron (TeV) in shot setup, Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting on the new store, and with the MiniBooNE and NuMI experiments taking beam.

At 8:18 AM, Operations reported that NuMI was off due to a kicker magnet flow fault.

At 8:49 AM, Operations reported having problems switching Pbar to its shot lattice. Operators called in experts. An expert discovered that there were some issues with the VSA code. He fixed the code.

A NuMI expert reported at 9:48 AM that a kicker magnet had tripped off. He said that he needed an access into the tunnel to check the kicker cooling fluid level.

Operations established combination store 4397 at 11:03 AM with an initial luminosity of 119.19E30. Pbar resumed stacking.

Operators began preparing the Main Injector (MI) and Booster at noon for an access.

At 12:16 PM, Operations reported that the QPMs and BPMs at TeV sectors A3 and A4 weren't responding. An expert investigated and found that a diagnostic device plugged into the network had caused the problem. They disconnected the device.

A Booster expert investigated the Booster notcher at 12:20 PM. Later he reported that he had activated an inhibit feature to stop the notcher overvoltage trips.

Operations reestablished beam in Linac at 2:59 PM.

NuMI experts reported at 3:09 PM that their kicker magnet cooling was back to normal.

Booster came out of access at 3:30 PM.

Operations established beam to the Booster dump at 4:17 PM.

An RF expert reported at 5:09 PM that all the MI RF stations were up and running.

MI came out of access and Operations began ramping the MI magnets at 5:26 PM.

Operations established beam to NuMI at 5:36 PM.

Pbar resumed stacking at 5:53 PM.

Operations prepared the Recycler to stash antiprotons at 7:59 PM. The stash ended and at 9:01 PM. Pbar resumed stacking.

Thursday September 22
The midnight shift began with Operations monitoring store 4397 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing for a stash at 12:48 AM. The stash ended at 1:41 AM. Pbar resumed stacking.

At 2:29 AM, a lightning strike caused a MI65 electrical safety trip and a NuMI trip as well. Operators investigated and found no real indication of system failures.

At 5:01 AM, Operations reported a TeV sector F0 vacuum burst. Operators reset the vacuum pumps.

The Recycler began preparing to stash at 5:19 AM. The stash ended and Pbar resumed stacking at 6:09 AM.

MCenter experimenters reported that they were preparing to install equipment and asked to have their beam disabled.

At 8:04 AM, Operations reported that Pbar RF cavities 1 and 8 were not putting out any voltages. Operations contacted an expert.

At 8:20 AM, Operations reported that a Debuncher RF station (DRF1) wasn't pulsing. Experts investigated and found a bad connection.

TeV experts began an end-of-store study at 1 PM.

Operations terminated store 4397 at 2:30 PM.

Operations put the TeV into shot setup at 4:20 PM.

Operations established combination store 4399 at 6:47 PM with an initial luminosity of 109.99E30. Pbar resumed stacking.

At 7:44 PM, Operations reported that stacking was poor due to DRF1. Experts investigated but couldn't figure out the cause of the problem without more investigations except that it had to do with phase jump circuitry, so they decided to switch to the spare board. Stacking improved at 9:53 PM.

Recycler began preparing to stash around 11:40 PM.

Friday September 23
The midnight shift began with Operations monitoring store 4399 and the stash, and with MiniBooNE and NuMI taking beam.

The Recycler stash ended and Pbar resumed stacking at 12:43 AM.

The Recycler began preparing to stash at 4:19 AM and ended stashing at 5:45 AM.

At 6:03 AM, Operations reported that they were having trouble getting Pbar back into stacking mode. Operations called in experts. An operator later discovered and resolved a DRF1 timing problem.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store. They also include Pbar and MI studies this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/23/2005   email Fermilab